Solve the system:
-x-y-2z=9
-2x-2y-z=1
-x-y+z=-10

Answer by ikleyn(52781)

.
- x - y -2z =   9,     (1)
-2x -2y - z =   1,     (2)
- x - y + z = -10.     (3)


Distract equation  (1)  from  equatin (3).  You will get

3z = -10 -9 = -19.     (4)


Now multiply equation  (3)  by  2  and then distract the equation  (2).  You will get

3z = 2*(-10) - 1 = -20 -1 = -21.     (5)


You have two equations for  z,  (4)  and  (5),  and they are inconsistent.


Hence,  the original system is inconsistent.


Answer.  This system has no solutions.



By the way,  the determinant of the system is equal to zero.